# Oscillating kink behavior in a traversable wormhole

### v1

@@ I Introduction
In field theories with spontaneous symmetry breaking, the causal disconnection between different regions of the system allows each region to select a different vacuum state. Topological defects then naturally emerge at the boundaries between these regions [brandenberger1994topological, teo2017topological, kosterlitz2017nobel]. Depending on the class of symmetry being broken, different classes of defects, such as monopoles, cosmic strings, and domain walls may emerge [pismen1999vortices, manton2004topological, bunkov2000topological]. Domain walls (also commonly referred to as kinks) arise from the spontaneous breaking of $Z_{2}$ symmetry. They exhibit a step-like structure in which the scalar field undergoes a sharp transition from one vacuum to the other, leading to a localized concentration of energy density at the interface. Owing to their extended nature and relatively high tension, domain walls can induce significant gravitational effects and play an important role in early-universe cosmology, phase transitions, and nonperturbative field dynamics [kibble1982phase, vilenkin1994cosmic, vachaspati2006kinks]. Domain walls are also often considered potential candidates for dark matter or other intriguing structures [press1989dynamical, avelino2008dynamics, friedland2003domain]. Fundamental constants, such as particle masses and the fine-structure constant and etc., may vary upon crossing a domain wall [kibble1976topology, vilenkin1982cosmic, vilenkin1985cosmic]. Since domain walls can alter fundamental physical constants, scientists have developed various experiments to detect them, including satellite synchronization [derevianko2014hunting, roberts2017search, kalaydzhyan2017extracting], electric dipole moment measurements [stadnik2014searching], magnetometry [pustelny2013global, afach2021search], acceleration induced by mass differences [mcnally2020constraining], and gravitational wave detectors [hall2018laser, grote2019novel, jaeckel2021probing]. Therefore, studying domain walls or kinks in gravitational systems is a task of considerable importance.
@@
The evolution of domain walls in flat spacetime has been extensively studied, including in both $2+1$ and $3+1$ dimensions. Reference [kevrekidis2018planar] examines the dynamical evolution of both planar and radial kinks in Klein–Gordon models, with a particular focus on their existence and stability properties. The attractive interaction between a kink and an antikink has been analyzed in Ref. [carretero2022kink], where an effective model was constructed via the variational approximation, yielding results in agreement with those from the full dynamical model. In Ref. [caputo2013radial], the influence of radial perturbations on kink configurations is studied, and the possibility of extracting energy from the kink is explored. More recently, some studies have employed the gauge/gravity duality [maldacena1999large, witten1998anti] to probe the statistical behavior of kinks in the strongly coupled regime of the boundary field theory [li2023black, ma2025universal]. Additional related work can be found in references [arodz1998expansion, dobrowolski2008construction, dobrowolski2009kink, gatlik2021modeling, gorria2004kink, shi2024topologically].
@@
In this context, an interesting question arises: compared with flat spacetime, what characteristic features does the kink exhibit in curved spacetime? Compact astrophysical objects, including black holes [ruffini1971introducing, frolov2012black], neutron stars [lattimer2004physics, lattimer2014neutron], boson stars [jetzer1992boson, schunck2003general], and wormholes [hawking1988wormholes, morris1988wormholes], serve as natural laboratories for studying the nonlinear behavior of fields under strong gravity. Compared with flat spacetime, the behavior of domain walls in curved geometry is significantly richer due to the background curvature, horizons, and the diverse structures of spacetime topology. In black hole backgrounds, scientists have conducted several studies on domain walls, including their stationary and transient behaviors. Studies have shown that black holes exhibit a repulsive effect on static thick domain walls [moderski2003thick, morisawa2003thick, moderski2004reissner, moderski2006thick]. In addition, when a kink passes near a black hole, it displays behavior resembling ringdown modes [ficek2018planar]. Recently, scientists have extended the study of kink dynamics to compact stars, such as neutron stars and boson stars. They found that compact stars can slow down the motion of kinks and may even produce a rebound-like effect on them [caputo2024radial, ma2025radial].
@@
However, studies of kinks in wormhole backgrounds remain very limited, with current research mostly focusing on static configurations and perturbative analyses [sushkov2002wormholes, bizon2021sine, diaz2023kinks, waterhouse2019phi, goulart2017phantom]. A wormhole refers to a tunnel-like spacetime structure that links two distant or otherwise disconnected regions. The earliest wormhole solution, known as the Einstein–Rosen bridge, was proposed by Einstein and Rosen [einstein1935particle]. Since this type of wormhole is non-traversable, it was long considered merely a mathematical construction rather than a physically realizable entity. Many years later, Ellis obtained a new wormhole solution, and Morris and Thorne subsequently proved that the Ellis wormhole is traversable, such that matter or light signals can safely pass through the wormhole without encountering horizons or singularities [ellis1973ether, morris1988wormholes2]. The existence of traversable wormholes requires matter that violates classical energy conditions, known as “exotic matter”, whose negative energy density can support the wormhole throat against collapse [chianese2017characterising].
@@ II.1 Simpson-Visser spacetime and wormhole configurations
The Simpson-Visser spacetime can be described by the line element presented in [simpson2019black],
@@ II.2 Radial kink model
which has minimums at $\phi=\pm 1$. Therefore, after breaking the $Z_{2}$ symmetry in the action (6), the scalar field $\phi$ will settle down to $\phi=+1$ or $\phi=-1$, which will in turn form the solutions of kinks [kibble1976topology, kibble1980some, zurek1985cosmological, ma2024asymmetric]. The equation of motion for the scalar field $\phi$ reads
@@ III Radial kinks in the Simpson-Visser wormholes
In this section we will study the evolution of kinks in the background of the Simpson-Visser wormholes. The dynamics of kinks are governed by the Eq. (II.2). The initial configuration of the kink can be given by [Weinberg:1992hc],
@@ IV Conclusions and Discussions
Due to the fact that the topology of the wormhole differs from that of ordinary compact stars, a single radial kink is protected by the wormhole topology and does not disappear at late times. Each time the radial kink passes through the wormhole throat, it transfers energy to the background in the form of emitted wave packets, thereby reducing its own energy and causing its oscillation amplitude to gradually decrease. This only drives the kink to oscillate within a progressively smaller region near the throat, while the kink itself persists throughout the evolution. In contrast, in the backgrounds of boson stars and neutron stars, the kink ultimately transfers all of its energy to the background through oscillations, and eventually the kink disappears into the background [caputo2024radial, ma2025radial].

### v2

@@ I Introduction
In field theories with spontaneous symmetry breaking, the causal disconnection between different regions of the system allows each region to select a different vacuum state. Topological defects then naturally emerge at the boundaries between these regions [7, 72, 46]. Depending on the class of symmetry being broken, different classes of defects, such as monopoles, cosmic strings, and domain walls may emerge [62, 54, 8]. Domain walls (also commonly referred to as kinks) arise from the spontaneous breaking of $Z_{2}$ symmetry. They exhibit a step-like structure in which the scalar field undergoes a sharp transition from one vacuum to the other, leading to a localized concentration of energy density at the interface. Owing to their extended nature and relatively high tension, domain walls can induce significant gravitational effects and play an important role in early-universe cosmology, phase transitions, and nonperturbative field dynamics [43, 75, 73]. Domain walls are also often considered potential candidates for dark matter or other intriguing structures [63, 3, 31]. Fundamental constants, such as particle masses and the fine-structure constant and etc., may vary upon crossing a domain wall [44, 74, 76]. Since domain walls can alter fundamental physical constants, scientists have developed various experiments to detect them, including satellite synchronization [23, 65, 41, 15], electric dipole moment measurements [70], magnetometry [64, 1], acceleration induced by mass differences [55], and gravitational wave detectors [37, 36, 39]. Therefore, studying domain walls or kinks in gravitational systems is a task of considerable importance.
@@
The evolution of domain walls in flat spacetime has been extensively studied, including in both $2+1$ and $3+1$ dimensions. Reference [42] examines the dynamical evolution of both planar and radial kinks in Klein–Gordon models, with a particular focus on their existence and stability properties. The attractive interaction between a kink and an antikink has been analyzed in Ref. [11], where an effective model was constructed via the variational approximation, yielding results in agreement with those from the full dynamical model. In Ref. [9], the influence of radial perturbations on kink configurations is studied, and the possibility of extracting energy from the kink is explored. More recently, some studies have employed the gauge/gravity duality [53, 79] to probe the statistical behavior of kinks in the strongly coupled regime of the boundary field theory [49, 50]. Additional related work can be found in references [2, 26, 27, 33, 34, 68].
@@
In this context, an interesting question arises: compared with flat spacetime, what characteristic features does the kink exhibit in curved spacetime? Compact astrophysical objects, including black holes [66, 32], neutron stars [47, 48], boson stars [40, 67], and wormholes [38, 61, 14], serve as natural laboratories for studying the nonlinear behavior of fields under strong gravity. Compared with flat spacetime, the behavior of domain walls in curved geometry is significantly richer due to the background curvature, horizons, and the diverse structures of spacetime topology. In black hole backgrounds, scientists have conducted several studies on domain walls, including their stationary and transient behaviors. Studies have shown that black holes exhibit a repulsive effect on static thick domain walls [56, 59, 57, 58]. In addition, when a kink passes near a black hole, it displays behavior resembling ringdown modes [30]. Recently, scientists have extended the study of kink dynamics to compact stars, such as neutron stars and boson stars. They found that compact stars can slow down the motion of kinks and may even produce a rebound-like effect on them [10, 52].
@@
However, studies of kinks in wormhole backgrounds remain very limited, with current research mostly focusing on static configurations and perturbative analyses [71, 6, 25, 77, 35]. A wormhole refers to a tunnel-like spacetime structure that links two distant or otherwise disconnected regions. The earliest wormhole solution, known as the Einstein–Rosen bridge, was proposed by Einstein and Rosen [28]. Since this type of wormhole is non-traversable, it was long considered merely a mathematical construction rather than a physically realizable entity. Many years later, Ellis obtained a new wormhole solution, and Morris and Thorne subsequently proved that the Ellis wormhole is traversable, such that matter or light signals can safely pass through the wormhole without encountering horizons or singularities [29, 60]. The existence of traversable wormholes requires matter that violates classical energy conditions, known as “exotic matter”, whose negative energy density can support the wormhole throat against collapse [12, 24]. For further studies on wormholes, see Refs. [5, 13, 18, 19, 17, 16, 4, 22, 21, 20].
@@ II.1 Simpson-Visser spacetime and wormhole configurations
The Simpson-Visser spacetime can be described by the line element presented in [69],
@@ II.2 Radial kink model
which has minimums at $\phi=\pm 1$. Therefore, after breaking the $Z_{2}$ symmetry in the action (6), the scalar field $\phi$ will settle down to $\phi=+1$ or $\phi=-1$, which will in turn form the solutions of kinks [44, 45, 80, 51]. The equation of motion for the scalar field $\phi$ reads
@@ III Radial kinks in the Simpson-Visser wormholes
In this section we will study the evolution of kinks in the background of the Simpson-Visser wormholes. The dynamics of kinks are governed by the Eq. (II.2). The initial configuration of the kink can be given by [78],
@@ IV Conclusions and Discussions
Due to the fact that the topology of the wormhole differs from that of ordinary compact stars, a single radial kink is protected by the wormhole topology and does not disappear at late times. Each time the radial kink passes through the wormhole throat, it transfers energy to the background in the form of emitted wave packets, thereby reducing its own energy and causing its oscillation amplitude to gradually decrease. This only drives the kink to oscillate within a progressively smaller region near the throat, while the kink itself persists throughout the evolution. In contrast, in the backgrounds of boson stars and neutron stars, the kink ultimately transfers all of its energy to the background through oscillations, and eventually the kink disappears into the background [10, 52].